Merrill Lynch, a wealth management and financial services company, is considering recommending a new stock to its clients. Prior to recommending it, they want to be able to say that the stock has outperformed the market return of 7% in the last ten years. To be as detailed as possible, they use monthly return data for the past 3 years. The average return for the stock is 73% and the known population standard deviation for stocks is 3%. Note, use whole numbers for percent (example for 73% plugin 7.3 not .073) D Calculate the appropriate test statistic needed to test your claim. Round your final answer to 4 decimals. Hint: Make sure you are using the correct formula based on the information given in the problem
Answer #1

Sol:

H0:\mu =7%

Ha:\mu >7%

3years

n=3*12=36 months

test statistic

z=sample mean-population mean/population stddev/sqrt(n)

=7.3-7/3/sqrt(36)

z=0.6

ANSWER:

0.6